# Resummed Distribution Functions: Making Perturbation Theory Positive and Normalized

### v1

@@ 4.1 Analytic matching
In Figs. 1 and 2, we show unitary completions of the exponential distribution and Rayleigh distribution, as given by Eq. (31) and Eq. (33), for random choices of the higher-order $g_{m}(t)$ parameters as described above. Each thin line is a different choice of the higher-order terms of $g_{M}(t)$, with the color representing the order $M$ of the approximation. Every single line here is positive, finite, and normalized, and it matches the perturbative expansion of the exponential up to and including $\mathcal{O}(\alpha^{M})$. When the higher-order terms are chosen to be zero, the RDF exactly reproduces the truth distribution. We first note that, at least by eye, all possible choices track the true distribution, which indicates that this “prior” on $g$ is reasonable. Second, we note that the distributions tighten as $M$ increases. This reflects our greater degree of perturbative certainty; in the $M\to\infty$ limit, the FO expansion is exactly equal to the true distribution, and there is no extra freedom that can be accessed by varying higher-order coefficients. [^13] The tightening is not merely due to the factor of $\frac{1}{m!}$ in the prior — it persists even when this is removed.
@@ 5.2 Multivariate Jet Angularities
In Fig. 6, we show the RDF multi-angularity given by Eq. (50) as well as the usual higher-order random variations. Here, we only vary $g_{\alpha}$ and not $g_{\beta}$, as the latter is subleading. The variations are the same as those in Sec. 5.1, though now we have 2D polynomials rather than 1D. This figure is the higher-dimensional analogue of Fig. 5 (more precisely, Fig. 5 is the marginal over $t_{\alpha}$ of Fig. 6). The RDF is able to properly normalize the fixed-order distribution while still respecting all kinematic boundaries. For comparison, the full LL result (Eq. (51)) is also shown. The LL result is captured by variations of $g_{\alpha}$, as expected, and visually, the LL result and the RDF result are similar.
@@
While we get the expected leading-log form for the $t_{\beta}$ marginal, we do not for the $t_{\alpha}$ marginal (though both at least have the correct fixed-order $\alpha_{s}\to 0$ limit). One should not lose hope, however, as the usual Sudakov factor is hiding inside – it is just obscured by the presence of higher-order $g$ terms that are allowed to be ignored to the order we are working in. By expanding the error functions, we can derive the $g$-function corresponding to $q(t_{\alpha}|\alpha_{s})$:
@@ 5.3 Thrust: Numeric Matching up to $\mathcal{O}(\alpha_{s}^{3})$
In this section, we apply the numeric matching procedure from Sec. 2.4 to event thrust Brandt:1964sa ; Farhi:1977sg , to obtain RDF predictions up to $\mathcal{O}(\alpha_{s}^{3})$. To accomplish this, we will use EERAD3 Aveleira:2025svg to obtain numeric estimates of the fixed-order calculation before applying numeric RDF matching as was done for the toys in Sec. 4.2.
@@
where $i$ indexes particles, $\vec{p_{i}}$ is the 3-momentum of that particle in the center-of-mass frame, and $\vec{n}$ is the “thrust axis”. The thrust $\tau$ takes on values between 0, for perfectly back-to-back events, to 0.5, for perfectly isotropic events. We elect to work with $x=2\tau$ and $t=\log(1/2\tau)$ when interfacing with the RDF, as it is simplest when the dynamic range of the random variable is between 0 to 1 (in $x$ space) or 0 to $\infty$ (in $t$) space. At any finite order in perturbation theory, however, the dynamic range of thrust is limited due to kinematics: in particular, at leading order, $\tau$ only ranges from 0 to $\frac{1}{3}$.
@@
We then generate the distributions of the thrust $\tau=1-T$ using EERAD3v2’s makedist command. For a fixed value of $\alpha_{s}(m_{Z})$, we may generate histograms for $\log(\tau)$. Histograms are generated such that $\log(\tau)$ is between -10 and 0, with 200 uniform bins. The kinematically-disallowed values of the thrust histogram have bin values of zero. We repeat this for 320 values of $\alpha_{s}(m_{Z})$, uniformly chosen between 0.005 and 0.325. EERAD3v2 also reports a Monte-Carlo uncertainty and a “theory-uncertainty” (in quotations, since we will define a new theory uncertainty in Sec. 6) due to scale variations on each bin. We include both in quadrature as the total error used in the MSE loss Eq. (20). By default, the error reported on the kinematically-disallowed bins is zero. Rather than removing these bins entirely (as it is important to reproduce that these bins are empty), we “clip” all errors from below, such that zero-errors are replaced by the the error on the first nonzero bin. We emphasize that this choice is just as a numeric regulator to avoid divide-by-zero errors in Eq. (20) — in principle, if one could perfectly guess the form of all $\Theta$-functions in the ansatz without gradient descent, this regularization would not be necessary as the loss would be finite.
In Fig. 7, we show the results of numerically matching the event thrust at first, second, and third orders in $\alpha_{s}$. All training hyperparameters are given in Table 2. For comparison, we also show the thrust as generated with Pythia 8.3 Sjostrand:2014zea in $e^{+}e^{-}\to q\bar{q}$ events at the $Z$-pole. For all three orders, the Taylor-expanded ansatz sits almost perfectly on top of the EERAD3v2 target. However, there is a notable discrepancy between the $\alpha=0.12$ RDF ansatzes and the Pythia prediction at the default tune. [^19] The default Monash 2013 tune of Pythia Skands:2014pea uses $\alpha_{s}=0.1365$ as its FSR showering parameter. This differs from the expected $0.118$, but the Pythia value is based on the CMW Catani:1990rr scheme rather than $\overline{\text{MS}}$, and is also different due to tuning. In particular, at order 1, the RDF ansatz tends to underfill the thrust phase space: the maximum thrust is $\sim 0.35$ compared to the Pythia maximum of 0.4. At orders 2 and 3, the RDF correctly drops off quickly near this kinematic threshold, though it is not sharp because of the $T$ parameters in the numeric RDF. However, we should not expect Pythia and the RDF predictions to completely agree: the former is an LL calculation including additional effects such as hadronization (which is known to shift the distribution), kinematic conservation, etc, while the latter is simply a unitary extension of a fixed order calculation.
@@ 6.1 Fitting procedure
We assume Gaussian likelihoods for both the observable bin counts and the nuisance parameters, which is typically a good approximation in the limit of large statistics. In principle, one could use the full event likelihood for the first term ($\sum_{i\in\text{Data}}\log(q(t_{i}|\alpha_{i},\phi,\nu))$ since $q$ is by construction a probability distribution, though collider data is often not made public in this format. Likewise, the prior on parameters need not be Gaussian; this is simply a choice. [^22] We also tried lognormal priors with a variety of means and standard deviations, but they provided very similar results to the Gaussian priors. We take $\mu_{k}=0$ and $\sigma_{k}=\frac{\sigma}{m!n!}$ for the two $g$ matrices, where $\sigma$ is some $\mathcal{O}(1)$ scale that we are free to choose. For the $T$ and $\theta$ parameters, we choose not to include a prior term (equivalent to choosing a flat unnormalized prior). We do not claim that these particular prior choices are optimal, and we will discuss priors more thoroughly in the following subsection. We stress that the MSE is taken between the non-Taylor expanded RDF and data. This is as opposed to in the numeric matching procedure to determine $\phi$, where we take an MSE between the Taylor-expanded RDF and a potentially non-unitary target.
@@ 6.2 Example: Fits to toys
To construct the “data” used in the log likelihood function from Eq. (59), we generate histograms from the all-orders expressions given in Eq. (23). We use 250,000 samples over 40 bins in $x$-space, ranging from 0 to 1, and assign statistical errors of $\sqrt{N_{i}}$ to each bin. In Fig. 8, we show the results of extracting $\alpha_{s}$ to first, second, and third orders on the exponential toy. In the left panel, we show the “experimental pseudodata” corresponding to an exponential distribution and the RDFs at orders $\mathcal{O}(\alpha_{s}^{1})$, $\mathcal{O}(\alpha_{s}^{2})$, and $\mathcal{O}(\alpha_{s}^{3})$ — that is to say, the RDFs matched to first, second, and third order in $\alpha_{s}$ from Sec. 4.2, then minimized over one higher order in the nuisance parameter. In the right panel, we show the profile likelihood as a function of $\alpha$ using a Gaussian prior with $\sigma=1.0$ for the nuisance parameters. The recovery of $\alpha_{s}$ is overall successful: the first, second, and third-order confidence intervals are all roughly centered around the true value of $\alpha_{s}$. Further, we see that the $\mathcal{O}(\alpha_{s}^{1})$ confidence interval completely envelopes the $\mathcal{O}(\alpha_{s}^{2})$ confidence interval, which completely envelopes the $\mathcal{O}(\alpha_{s}^{3})$ interval. These results are what we might expect: higher-order RDFs contain more information about the underlying distribution, and thus more closely resemble the target all-orders PDF. For lower-order calculations, the nuisance parameters have to do less “work” to match the target data (since they multiply a lower power of $\alpha_{s}$), resulting in many more choices of nuisance parameters giving a valid fit.
@@
We show analogous fits to the Rayleigh distribution to first, second, and third order in $\alpha$ in Fig. 9. For this example, while the $\mathcal{O}(\alpha_{s}^{2})$ and $\mathcal{O}(\alpha_{s}^{3})$ confidence intervals are roughly centered around the true value of $\alpha_{s}$, the $\mathcal{O}(\alpha_{s}^{1})$ confidence interval seems to miss the mark, though it still contains the target within $2\sigma$. Of course, if well-calibrated, the confidence intervals should only be expected to cover the true value only $68\%$ of the time. Similar to the exponential toy, lower-order RDFs have more valid choices for the nuisance parameters, resulting in wider confidence intervals. Further, the $\mathcal{O}(\alpha_{s}^{2})$ confidence interval does not fully envelope the $\mathcal{O}(\alpha_{s}^{3})$ interval, although the amount of overlap is high. We do see that the confidence intervals shrink as the order of the RDF increases. This shrinkage implies again that embedding more information into the RDF during the matching procedure means that the nuisance parameters can do less work fitting the RDF to the target and instead be more effective at constraining $\alpha_{s}$.
@@ 6.3 Extracting $\alpha_{s}$ With Fits to ALEPH Data
In Fig. 10, we show the results of extracting $\alpha$ to $\mathcal{O}(\alpha_{s}^{1})$, $\mathcal{O}(\alpha_{s}^{2})$, and $\mathcal{O}(\alpha_{s}^{3})$, with a Gaussian prior on parameters of $\sigma=1.0$: We also show the extracted $\alpha_{s}$ values in Table 1 for each order and for $\sigma=0.5,1.0$, and $2.0$, as well as the results from Ref. Becher:2008cf for comparison. in the left panel of Fig. 10, we show the RDFs along with the ALEPH data they are fit to, and in the right panel, we show the profile likelihood as a function of $\alpha$ using a Gaussian prior for the nuisance parameters with a standard deviation of 1.0.272727We show the fit results for other choices of the prior standard deviation in App. A.1
@@
Given the similar qualitative results to our Rayleigh toy example, it seems likely that the lowest-order RDF does not contain enough information about the target PDF in order for the nuisance parameter minimization to be completely effective. In the case of $\sigma=2.0$ for example, the $\mathcal{O}(\alpha_{s}^{1})$ fit fails to find a sensible minimum, and simply returns the largest $\alpha_{s}$ value we scanned over [^28] In principle, we could have gone further, but the $\alpha_{s}$ returned is regardless much larger than one would expect. In general, the $\mathcal{O}(\alpha_{s}^{1})$ fits are finicky, and we had to use an additional L-BFGS minimizer Liu1989OnTL after the gradient descent to avoid getting stuck in local minima., and for $\sigma=0.5$, the $\mathcal{O}(\alpha_{s}^{1})$ result is several standard deviations away from sensible values. On the other hand, the second-order and higher fits seem relatively robust, and also do not change much as the prior $\sigma$ changes. For users interested in using the RDF framework for parameter fits, we recommend to match at least up to second order before minimizing over nuisance parameters.
@@ 6.4 Brief comparison to other methods
where $Q$ denotes some physical scale related to the calculation. Of course, physical quantities should be independent of the renormalization scale. However, truncating the sum to a finite order in $\alpha$ will still be $\mu_{R}$-dependent, so we can use $\mu_{R}$ as a lever to quantify the uncertainty on the higher-order terms in the sum. The conventional approach for citing these truncation uncertainties is to quote the physical observable at $\mu_{R}=Q$ (i.e. the physical scale of the problem), then give an error envelope defined over the variation of the renormalization scale from Q/2,2Q{Q/2,2Q}. Despite the widespread use of this conventional approach, it has several drawbacks For one, the variation by a factor of 2 is arbitrary. For another, there is no real probabilistic or physical interpretation of the variation of $\mu_{R}$ (we cannot, for instance say that the untruncated cross section has a 95% chance of lying within the cited envelope), unless this is considered as a subjective Bayesian prior.
As noted by Ref. Tackmann:2024kci , scale variation uncertainties can be insufficient to parameterize the space of all possible higher-order terms — in a sense, they only capture the physics that we “already know”. Our method is truly physics-agnostic, as we simply parameterize the form of higher-order coefficients. Note that one must still choose a finite parameterization for $g$ (e.g. choosing $g$ to be polynomial in $t$), and therefore a bias in the functional form, but in principle a set of complete basis functions may be chosen. Unlike the approach of Ref. Tackmann:2024kci , which also advocates for treating higher-order terms as nuisance parameters but still works at the level of fixed-order expressions, the RDF method guarantees that all possible variations are physical for a differential cross section.
@@
Our approach is similar to Cacciari-Houdeau (and modifications) in that we use prior-dependent bounds on coefficients. Both our method and the Cacciari-Houdeau method do not rely on the renormalization scale. However, we make stronger (yet well-motivated) assumptions on the coefficients of the perturbative series $\sigma=\sum_{m=0}^{\infty}c_{m}\alpha^{m}$ (unitarity) which allow us to write down a closed-form expression for the coefficients — the RDF itself. Further, we allow for more user input on the coefficients of the power series: while the Cacciari-Houdeau approach makes claims at the level of $c_{m}$, our approach allows us to specify the form of $c_{m}$, for example as a polynomial in $t$. In the case where there is a strong argument that $c_{m}$ should take a particular functional form in $t$, then our method allows for more fine-grained determination of the associated coefficients.

### v2

@@ 4.1 Analytic matching
In Figs. 1 and 2, we show unitary completions of the exponential distribution and Rayleigh distribution, as given by Eq. (31) and Eq. (33), for random choices of the higher-order $g_{m}(t)$ parameters as described above. Each thin line is a different choice of the higher-order terms of $g_{M}(t)$, with the color representing the order $M$ of the approximation. Every single line here is positive, finite, and normalized, and it matches the perturbative expansion of the exponential up to and including $\mathcal{O}(\alpha^{M})$. When the higher-order terms are chosen to be zero, the RDF exactly reproduces the truth distribution. We first note that, at least by eye, all possible choices track the true distribution, which indicates that this “prior” on $g$ is reasonable. Second, we note that the distributions tighten as $M$ increases. This reflects our greater degree of perturbative certainty: in the $M\to\infty$ limit, the FO expansion is exactly equal to the true distribution, and there is no extra freedom that can be accessed by varying higher-order coefficients. [^13] The tightening is not merely due to the factor of $\frac{1}{m!}$ in the prior — it persists even when this is removed.
@@ 5.2 Multivariate Jet Angularities
In Fig. 6, we show the RDF multi-angularity given by Eq. (50) as well as the usual higher-order random variations. Here, we only vary $g_{\alpha}$ and not $g_{\beta}$, as the latter is subleading. The variations are the same as those in Sec. 5.1, though now we have 2D polynomials rather than 1D. This figure is the higher-dimensional analogue of Fig. 5 (more precisely, Fig. 5 is the marginal over $t_{\alpha}$ of Fig. 6). The RDF is able to properly normalize the fixed-order distribution while still respecting all kinematic boundaries. For comparison, the full LL result (Eq. (51)) is also shown. The LL result is captured by variations of $g_{\alpha}$, as expected. Visually, the LL result and the RDF result are similar.
@@
While we get the expected leading-log form for the $t_{\beta}$ marginal, we do not for the $t_{\alpha}$ marginal (though both at least have the correct fixed-order $\alpha_{s}\to 0$ limit). One should not lose hope, however, as the usual Sudakov factor is hiding inside — it is just obscured by the presence of higher-order $g$ terms that are allowed to be ignored to the order we are working in. By expanding the error functions, we can derive the $g$-function corresponding to $q(t_{\alpha}|\alpha_{s})$:
@@ 5.3 Thrust: Numeric Matching up to $\mathcal{O}(\alpha_{s}^{3})$
In this section, we apply the numeric matching procedure from Sec. 2.4 to event thrust Brandt:1964sa ; Farhi:1977sg to obtain RDF predictions up to $\mathcal{O}(\alpha_{s}^{3})$. To accomplish this, we will use EERAD3 Aveleira:2025svg to obtain numeric estimates of the fixed-order calculation before applying numeric RDF matching as was done for the toys in Sec. 4.2.
@@
where $i$ indexes particles, $\vec{p_{i}}$ is the 3-momentum of that particle in the center-of-mass frame, and $\vec{n}$ is the “thrust axis”. The thrust $\tau$ takes on values between 0, for perfectly back-to-back events, to 0.5, for perfectly isotropic events. We elect to work with $x=2\tau$ and $t=\log(1/2\tau)$ when interfacing with the RDF, as it is simplest when the dynamic range of the random variable is between 0 to 1 (in $x$ space) or 0 to $\infty$ (in $t$ space). At any finite order in perturbation theory, however, the dynamic range of thrust is limited due to kinematics: in particular, at leading order, $\tau$ only ranges from 0 to $\frac{1}{3}$.
@@
We then generate the distributions of the thrust $\tau=1-T$ using EERAD3v2’s makedist command. For a fixed value of $\alpha_{s}(m_{Z})$, we may generate histograms for $\log(\tau)$. Histograms are generated such that $\log(\tau)$ is between -10 and 0, with 200 uniform bins. The kinematically-disallowed values of the thrust histogram have bin values of zero. We repeat this for 320 values of $\alpha_{s}(m_{Z})$, uniformly chosen between 0.005 and 0.325. EERAD3v2 also reports a Monte-Carlo uncertainty and a “theory-uncertainty” (in quotations, since we will define a new theory uncertainty in Sec. 6) due to scale variations on each bin. We include both in quadrature as the total error used in the MSE loss Eq. (20). By default, the error reported on the kinematically-disallowed bins is zero. Rather than removing these bins entirely (as it is important to reproduce that these bins are empty), we “clip” all errors from below, such that zero-errors are replaced by the error on the first nonzero bin. We emphasize that this choice is just as a numeric regulator to avoid divide-by-zero errors in Eq. (20) — in principle, if one could perfectly guess the form of all $\Theta$-functions in the ansatz without gradient descent, this regularization would not be necessary as the loss would be finite.
In Fig. 7, we show the results of numerically matching the event thrust at first, second, and third orders in $\alpha_{s}$. All training hyperparameters are given in Table 2. For comparison, we also show the thrust as generated with Pythia 8.3 Sjostrand:2014zea in $e^{+}e^{-}\to q\bar{q}$ events at the $Z$-pole. For all three orders, the Taylor-expanded ansatz sits almost perfectly on top of the EERAD3v2 target. However, there is a notable discrepancy between the $\alpha=0.12$ RDF ansatzes and the Pythia prediction at the default tune. [^19] The default Monash 2013 tune of Pythia Skands:2014pea uses $\alpha_{s}=0.1365$ as its FSR showering parameter. This differs from the expected $0.118$, but the Pythia value is based on the CMW Catani:1990rr scheme rather than $\overline{\text{MS}}$, and is also different due to tuning. In particular, at order 1, the RDF ansatz tends to underfill the thrust phase space: the maximum thrust is $\sim 0.35$ compared to the Pythia maximum of 0.4. At orders 2 and 3, the RDF correctly drops off quickly near this kinematic threshold, though it is not sharp because of the $T$ parameters in the numeric RDF. However, we should not expect the Pythia and the RDF predictions to completely agree: the former is an LL calculation including additional effects such as hadronization (which is known to shift the distribution), kinematic conservation, etc, while the latter is simply a unitary extension of a fixed order calculation.
@@ 6.1 Fitting procedure
We assume Gaussian likelihoods for both the observable bin counts and the nuisance parameters, which is typically a good approximation in the limit of large statistics. In principle, one could use the full event likelihood for the first term ($\sum_{i\in\text{Data}}\log(q(t_{i}|\alpha_{i},\phi,\nu))$) since $q$ is by construction a probability distribution, though collider data is often not made public in this format. Likewise, the prior on parameters need not be Gaussian; this is simply a choice. [^22] We also tried lognormal priors with a variety of means and standard deviations, but they provided very similar results to the Gaussian priors. We take $\mu_{k}=0$ and $\sigma_{k}=\frac{\sigma}{m!n!}$ for the two $g$ matrices, where $\sigma$ is some $\mathcal{O}(1)$ scale that we are free to choose. For the $T$ and $\theta$ parameters, we choose not to include a prior term (equivalent to choosing a flat unnormalized prior). We do not claim that these particular prior choices are optimal, and we will discuss priors more thoroughly in the following subsection. We stress that the MSE is taken between the non-Taylor expanded RDF and data. This is as opposed to in the numeric matching procedure to determine $\phi$, where we take an MSE between the Taylor-expanded RDF and a potentially non-unitary target.
@@ 6.2 Example: Fits to toys
To construct the “data” used in the log likelihood function from Eq. (59), we generate histograms from the all-orders expressions given in Eq. (23). We use 250,000 samples over 40 bins in $x$-space, ranging from 0 to 1, and we assign statistical errors of $\sqrt{N_{i}}$ to each bin. In Fig. 8, we show the results of extracting $\alpha_{s}$ to first, second, and third orders on the exponential toy. In the left panel, we show the “experimental pseudodata” corresponding to an exponential distribution and the RDFs at orders $\mathcal{O}(\alpha_{s}^{1})$, $\mathcal{O}(\alpha_{s}^{2})$, and $\mathcal{O}(\alpha_{s}^{3})$ — that is to say, the RDFs matched to first, second, and third order in $\alpha_{s}$ from Sec. 4.2, then minimized over one higher order in the nuisance parameter. In the right panel, we show the profile likelihood as a function of $\alpha$ using a Gaussian prior with $\sigma=1.0$ for the nuisance parameters. The recovery of $\alpha_{s}$ is overall successful: the first, second, and third-order confidence intervals are all roughly centered around the true value of $\alpha_{s}$. Further, we see that the $\mathcal{O}(\alpha_{s}^{1})$ confidence interval completely envelopes the $\mathcal{O}(\alpha_{s}^{2})$ confidence interval, which completely envelopes the $\mathcal{O}(\alpha_{s}^{3})$ interval. These results are what we might expect: higher-order RDFs contain more information about the underlying distribution, and thus more closely resemble the target all-orders PDF. For lower-order calculations, the nuisance parameters have to do less “work” to match the target data (since they multiply a lower power of $\alpha_{s}$), resulting in many more choices of nuisance parameters giving a valid fit.
@@
We show analogous fits to the Rayleigh distribution to first, second, and third order in $\alpha$ in Fig. 9. For this example, while the $\mathcal{O}(\alpha_{s}^{2})$ and $\mathcal{O}(\alpha_{s}^{3})$ confidence intervals are roughly centered around the true value of $\alpha_{s}$, the $\mathcal{O}(\alpha_{s}^{1})$ confidence interval seems to miss the mark, though it still contains the target within $2\sigma$. Of course, if well-calibrated, the confidence intervals should only be expected to cover the true value $68\%$ of the time. Similar to the exponential toy, lower-order RDFs have more valid choices for the nuisance parameters, resulting in wider confidence intervals. Further, the $\mathcal{O}(\alpha_{s}^{2})$ confidence interval does not fully envelope the $\mathcal{O}(\alpha_{s}^{3})$ interval, although the amount of overlap is high. We do see that the confidence intervals shrink as the order of the RDF increases. This shrinkage implies again that embedding more information into the RDF during the matching procedure means that the nuisance parameters can do less work fitting the RDF to the target and instead be more effective at constraining $\alpha_{s}$.
@@ 6.3 Extracting $\alpha_{s}$ With Fits to ALEPH Data
In Fig. 10, we show the results of extracting $\alpha$ to $\mathcal{O}(\alpha_{s}^{1})$, $\mathcal{O}(\alpha_{s}^{2})$, and $\mathcal{O}(\alpha_{s}^{3})$, with a Gaussian prior on parameters of $\sigma=1.0$. We also show the extracted $\alpha_{s}$ values in Table 1 for each order and for $\sigma=0.5,1.0$, and $2.0$, as well as the results from Ref. Becher:2008cf for comparison. In the left panel of Fig. 10, we show the RDFs along with the ALEPH data they are fit to, and in the right panel, we show the profile likelihood as a function of $\alpha$ using a Gaussian prior for the nuisance parameters with a standard deviation of 1.0.272727We show the fit results for other choices of the prior standard deviation in App. A.1
@@
Given the similar qualitative results to our Rayleigh toy example, it seems likely that the lowest-order RDF does not contain enough information about the target PDF in order for the nuisance parameter minimization to be completely effective. In the case of $\sigma=2.0$ for example, the $\mathcal{O}(\alpha_{s}^{1})$ fit fails to find a sensible minimum, and simply returns the largest $\alpha_{s}$ value we scanned over [^28] In principle, we could have gone further, but the $\alpha_{s}$ returned is regardless much larger than one would expect. In general, the $\mathcal{O}(\alpha_{s}^{1})$ fits are finicky, and we had to use an additional L-BFGS minimizer Liu1989OnTL after the gradient descent to avoid getting stuck in local minima., and for $\sigma=0.5$, the $\mathcal{O}(\alpha_{s}^{1})$ result is several standard deviations away from sensible values. On the other hand, the second-order and higher fits seem relatively robust, and they also do not change much as the prior $\sigma$ changes. For users interested in using the RDF framework for parameter fits, we recommend to match at least up to second order before minimizing over nuisance parameters.
@@ 6.4 Brief comparison to other methods
where $Q$ denotes some physical scale related to the calculation. Of course, physical quantities should be independent of the renormalization scale. However, truncating the sum to a finite order in $\alpha$ will still be $\mu_{R}$-dependent, so we can use $\mu_{R}$ as a lever to quantify the uncertainty on the higher-order terms in the sum. The conventional approach for citing these truncation uncertainties is to quote the physical observable at $\mu_{R}=Q$ (i.e. the physical scale of the problem), then give an error envelope defined over the variation of the renormalization scale from Q/2,2Q{Q/2,2Q}. Despite the widespread use of this conventional approach, it has several drawbacks. For one, the variation by a factor of 2 is arbitrary. For another, there is no real probabilistic or physical interpretation of the variation of $\mu_{R}$ (we cannot, for instance say, that the untruncated cross section has a 95% chance of lying within the cited envelope), unless this is considered as a subjective Bayesian prior.
As noted by Ref. Tackmann:2024kci , scale variation uncertainties can be insufficient to parameterize the space of all possible higher-order terms — in a sense, they only capture the physics that we “already know”. Our method is truly physics-agnostic, as we simply parameterize the form of higher-order coefficients. Note that one must still choose a finite parameterization for $g$ (e.g. choosing $g$ to be polynomial in $t$), and therefore a bias in the functional form, but in principle a set of complete basis functions may be chosen. Unlike the approach of Ref. Tackmann:2024kci , which also advocates for treating higher-order terms as nuisance parameters, the RDF method guarantees that all possible variations are physical for a differential cross section. An interesting difference to point out between the Theory Nuisance Parameter (TNP) approach in resummed calculations and our method is that while the perturbative objects $f_{n}$ in the TNP approach (corresponding to our $p_{m}$ in Eq. (1)) take advantage of physics structure (color factors, anomalous dimension structure, etc), the only structure from our RDF approach comes from enforcing unitarity. [^29] For a similar study extracting $\alpha_{s}$ from a resummed $Z$ boson $q_{T}$ spectrum using TNPs, see Ref. Cridge:2025wwo .
@@
Our approach is similar to Cacciari-Houdeau (and modifications) in that we use prior-dependent bounds on coefficients. Both our method and the Cacciari-Houdeau method do not rely on the renormalization scale. However, we make stronger (yet well-motivated) assumptions on the coefficients of the perturbative series $\sigma=\sum_{m=0}^{\infty}c_{m}\alpha^{m}$ (unitarity) which allow us to write down a closed-form expression for the coefficients — the RDF itself. Further, we allow for more user input on the coefficients of the power series: while the Cacciari-Houdeau approach makes claims at the level of $c_{m}$, our approach allows us to specify the form of $c_{m}$, for example as a polynomial in $t$. In the case where there is a strong argument that $c_{m}$ should take a particular functional form in $t$, our method allows for more fine-grained determination of the associated coefficients.